### v1

# Sensor Resistant Instruction Independent Obfuscation for Multiple Programs

Abstract.
This work builds upon and optimizes our prior research on obfuscation as instruction decorrelation which achieves multiple program obfuscation. Leveraging this infrastructure, we further achieve the property of sensor-resistant computation.
Software Protection, Obfuscation, Instruction Independence, Instruction Decorrelation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 2023 ACM SIGPLAN International Symposium on New Ideas, New Paradigms, and Reflections on Programming and Software; October 20–25, 2024; Pasadena, California, USAisbn: 978-1-4503-XXXX-X/18/06copyright: noneccs: Software and its engineering Software organization and properties
1. Introduction
Accessing the binaries of a computer program raises an important question of how much can we learn about the program beyond simply running it with different inputs and observing its behavior. There are numerous techniques available that allow adversaries to analyze these binaries or run the program to study its behavior and internal states, with the intention of reverse engineering the source code or extracting secrets stored within the program. To protect software from such attacks, computer science has employed various ad-hoc heuristics for nearly four decades (Xu et al., 2020) and formal models for over two decades (Barak et al., 2001; Goldwasser and Rothblum, 2007; Kuzurin et al., 2007; Ali Ajorian, 2024), collectively referred to as program obfuscation or shortly obfuscation. Roughly speaking, obfuscation is a transformation that maintains a program’s functionality while hindering the extraction of valuable information, such as algorithms, data structures, and secret keys, with an acceptable slowdown overhead.
Ajorian et. al introduced obfuscation as instruction decorrelation, which focuses on the internal structures of regular programs to hide essential instructions within a set of junk instructions (Ali Ajorian, 2024). They defined the unintelligibility of an obfuscator as the decorrelation of all possible pairs of essential and junk instructions, ensuring that an efficient adversary cannot identify a correlated set of instructions to reconstruct an obfuscated program.
In this work, we enhance the generation of junk instructions by obfuscating a collection of programs rather than focusing on a single program. Specially, for each program the remaining programs serve as the source of junk instructions. Since the obfuscator ensures that all obfuscated instructions are decorrelated, no efficient adversary can determine the origin of each instruction to disrupt the execution of a program. This means that an evaluator executing the output of such an obfuscator can run either all or none of the input programs, making it impossible to censor any of the input programs.
2. Preliminaries
This section introduces the notation that will be used consistently across this work and offers foundational background information relevant to our study.
2.1. Notation
PPT stands for probabilistic polynomial time algorithms, which can make random choices during execution, and their running time is bounded by a polynomial function of the input size.
By $x\leftarrow D$ we mean that $x$ is a random variable drawn from the distribution $D$. The notation $A(.)\rightarrow x$ denotes that a probabilistic algorithm A generates an outcome $x$ from an input (.)(.)( . ).
The notation |.||.|| . |, in addition to its standard mathematical meanings such as the length of a set, is also utilized for indicating program size, referring to the number of elements or instructions within the program.
If $X$ is a set, by $X^{*}$ we mean all possible sequences of any length defined on the set X.
A function $\epsilon:\mathbb{N}\rightarrow\mathbb{N}$ is defined as negligible function if its growth rate is slower than the inverse of any polynomial. In other words, for any positive polynomial $p(n)$, there exists a threshold $n_{0}$ such that for all $n>n_{0}$, $\epsilon(n)<\frac{1}{p(n)}$.
By $set(p)$, we mean the set of all instructions in the program $p$.
2.2. Background and Related Works
Let $\mathcal{P}$ be the set of all possible programs defined on an instruction space $\mathcal{I}$, Ajorian et. al defined an instruction-independent obfuscation (Ali Ajorian, 2024) as a PPT algorithm $IIO:\mathcal{P}\rightarrow\mathcal{P}$ with the following properties:
  • Functionality preservation: $\hat{p}=IIO(p)$ computes the same functionality as input program $p$.
  • Polynomial slowdown: There exist polynomials $q$ and $r$ such that $|IIO(p)|\leq|p|$ and if $p$ halts within a maximum $t$ steps on input $x$, the obfuscated program halts within $r(t)$ steps on the same input.
  • Unintelligibility: for any auxiliary input $z$ of polynomial size, for any PPT adversary $A$ there must exist a negligible function $\epsilon$ such that:
    (1) $\begin{split}\big{|}&\Pr\nolimits_{A(\hat{p},z)\rightarrow\hat{s}_{1},\hat{s}_{2};\hat{s}_{1}\neq\hat{s}_{2}}[\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{S}\text{ or }\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{J}]-\\ &\Pr\nolimits_{\hat{s}_{1},\hat{s}_{2}\leftarrow\Omega;\hat{s}_{1}\neq\hat{s}_{2}}[\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{S}\text{ or }\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{J}]\big{|}\leq\epsilon(\lambda)\end{split}$
where $\lambda$ is a user choice of a security parameter, $\mathbb{S}$ and $\mathbb{J}$ are the sets of essential and junk instructions respectively and $\Omega=set(\hat{p})$.
3. IIO with Multiple Programs
An IIO obfuscator employs a set of junk instructions, denoted as $\mathbb{J}$, with specific requirements to conceal the set of essential instructions $\mathbb{S}$ of a source program $p$. Provided these requirements are satisfied, $\mathbb{J}$ can be constructed using various methods. Specially, $\mathbb{J}$ may consist of the union of several $\mathbb{S}_{i}$s where each $\mathbb{S}_{i}$ represents the essential instructions of a program $p_{i}$. From this perspective, an IIO obfuscator can be viewed as a transformation $\mathcal{O}$ that takes a set of programs $P=\{p_{1},\ldots,p_{n}\}$ as inputs and produces a single obfuscated program $\mathcal{O}(P)$ that is semantically equivalent to the set $P$ while ensuring that an efficient adversary cannot correlate any pair of instructions originating from the same source program.
The following definition encapsulates this concept of obfuscation:
Definition 3.1 (Multiple Program IIO).
A PPT algorithm $\mathcal{O}$ is considered as a multiple program IIO for a set of input programs $P=\{p_{1},\ldots,p_{n}\}$ if its output satisfies the following three conditions:
  • Functionality preservation: $\mathcal{O}(P)$ computes the same functionality as running all individual programs in $P$.
  • Polynomial slowdown: There exist polynomials $q$ and $r$ such that $|\mathcal{O}(P)|\leq q(\sum_{i=1}^{n}|p_{i}|)$ and if $p_{1},\ldots,p_{n}$ halts within a maximum $t$ of steps on inputs $x_{1},\ldots,x_{n}$ respectively, the obfuscated program $\mathcal{O}(P)$ halts within $r(t)$ steps on the same input.
  • Unintelligibility: for any auxiliary input $z$ of polynomial size, for any PPT adversary $A$ and for all $i\in\{1,...,n\}$ there exist a negligible function $\epsilon$ such that:
    (2) $\begin{split}\big{|}&\Pr\nolimits_{A(\mathcal{O}(P),z)\rightarrow\hat{s}_{1},\hat{s}_{2};\hat{s}_{1}\neq\hat{s}_{2}}[\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{S}_{i}]-\\ &\Pr\nolimits_{\hat{s}_{1},\hat{s}_{2}\leftarrow\Omega;\hat{s}_{1}\neq\hat{s}_{2}}[\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{S}_{i}]\big{|}\leq\epsilon(\lambda)\end{split}$
    where $\lambda$ is a user choice of a security parameter and $\mathbb{S}_{i}$ is the set of essential programs for the program $p_{i}$ and $\mathbb{J}_{i}=\bigcup_{j=1,j\neq i}^{n}\mathbb{S}_{j}$.
Note that, what we have defined as unintelligibility in Equation 2 is not directly derived from Equation 1. Rather, Equation 3, presented below, is a direct rewriting of Equation 1 for multi-program obfuscation. Furthermore, in Theorem 3.2, we prove that Equations 2 and 3 are equivalent.
(3) $\begin{split}&\forall i\in\{1\ldots n\}:\\ \big{|}&\Pr\nolimits_{A(\mathcal{O}(P),z)\rightarrow\hat{s}_{1},\hat{s}_{2};\hat{s}_{1}\neq\hat{s}_{2}}[\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{S}_{i}\text{ or }\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{J}_{i}]-\\ &\Pr\nolimits_{\hat{s}_{1},\hat{s}_{2}\leftarrow\Omega;\hat{s}_{1}\neq\hat{s}_{2}}[\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{S}_{i}\text{ or }\{\hat{s}_{1},\hat{s}_{2}\}\subseteq\mathbb{J}_{i}]\big{|}\leq\epsilon(\lambda)\end{split}$
Theorem 3.2.
: Equation 2 and Equation 3 are equivalent.
Proof.
3.1. Resilience to Sensorship
4. Tampering Attack
Consider an obfuscated program $\mathcal{O}(P)$ that takes input vector $X=(x_{1},\ldots,x_{n})$ and produces output values $Y=(y_{1},\ldots,y_{n})$ for the programs $p_{1},\ldots,p_{n}$ respectively. Throughout this process, the obfuscated program encounters a sequence of internal states $S=(s_{1},\ldots,s_{m})$. An adversary with the capability of tampering with the instructions or the internal states of $\mathcal{O}(P)$ can monitors $X,Y$ and $S$ before and after the tampering. The differences in these observations could potentially provide the adversary some information to correlate instructions and violate security guarantee of Equation 2. Specially, the adversary could remove an instruction from the obfuscated program, execute the resulting program, and determine which output value $y_{i}$ changes. By doing this for polynomially many times, the adversary can detect correlations between the instructions of the obfuscated program.
4.1. Tamper-Resistance Output
To defend tampering attack, the obfuscator must guarantee that any attempt to tamper with $\mathcal{O}(P)$ either leaves the output vector $Y$ unchanged or or alters all its elements entirely. This can be accomplished through two approaches: (1) implementing a compensation mechanism that neutralizes the effects of tampered instructions, ensuring $Y$ remains unchanged, or (2) employing an encoding scheme for $Y$ that depends on all instructions of the obfuscated program, ensuring that any tampering affects the entire output vector uniformly.
5. Implementation
In this section, we present an implementation of an obfuscator that hides which source program an obfuscated instruction originated from (Theorem LABEL:th:singleinst) and data access dependencies between instructions (Theorem LABEL:th:data-access-decorrelation).
5.1. Architecture Overview
5.2. Adversarial Model
6. Security Analysis
7. Implementation Results
8. Conclusions and Open Problems
Acknowledgements.
We would like to express our sincere gratitude to Osman Bicer for numerous insightful discussions concerning the security properties of this work.